Wildhorse Leasing Company leases a new machine to Sharrer Corporation. The machine has a cost of $65,000 and fair value of $95,500. Under the 3-year, non-cancelable contract, Sharrer will receive title to the machine at the end of the lease. The machine has a 3-year useful life and no residual value. The lease was signed on January 1, 2020. Wildhorse expects to earn an 8% return on its investment, and this implicit rate is known by Sharrer. The annual rentals are payable on each December 31, beginning December 31, 2020.

Prepare an amortization schedule that would be suitable for both the lessor and the lessee and that covers all the years involved. (For calculation purposes, use 5 decimal places as displayed in the factor table provided and round final answers to 0 decimal places e.g. 5,275.)

Date

Rent Receipt/ Payment

Interest Revenue/ Expense

Reduction of Principal

Receivable/ Liability

1/1/20

$enter a dollar amount rounded to 0 decimal places

$enter a dollar amount rounded to 0 decimal places

$enter a dollar amount rounded to 0 decimal places

$enter a dollar amount rounded to 0 decimal places

12/31/20

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

12/31/21

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

12/31/22

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

enter a dollar amount rounded to 0 decimal places

Prepare the journal entry at commencement of the lease for Wildhorse. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

1/1/20

enter an account title for the journal entry on January 1

enter a debit amount

enter a credit amount

enter an account title for the journal entry on January 1

enter a debit amount

enter a credit amount

enter an account title for the journal entry on January 1

enter a debit amount

enter a credit amount

enter an account title for the journal entry on January 1

enter a debit amount

enter a credit amount

Prepare the journal entry at commencement of the lease for Sharrer. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

1/1/20

enter an account title for the journal entry on January 1

enter a debit amount

enter a credit amount

enter an account title for the journal entry on January 1

enter a debit amount

enter a credit amount
Answer #1

Fair value of Asset: 95500

Implicit discount rate = 8%

Term is 3 years. So, Lease rental = 95500/(PV 8% for 3 years)

ie., 95500/(2.5771) = 35057.

So,

Date Lease rent Interest Reduction of Principal Principal O/s
1/1/2020 95500
31/12/20 37057 95500*8/100 7640 37057-7640 = 29417 95500-29417 66083
31/12/21 37057 66083*8/100 5287 37057-5286 = 31771 66083-31771 34312
31/12/22 37057 34312*8/100 2745 37057-2745 = 34312 34312-34312 0

1.The journal entry at commencement of the lease for Wildhorse:

Lease rent receivable a/c Dr 95500
Cost of Goods sold a/c Dr 65000
To Sales Revenue 95500
To Inventory 65000

2. The journal entry at commencement of the lease for Sharrer:

Right of Use asset a/c Dr 95500
To Lease Liabiility a/c 95500